Que. 1 The milli moles of Borax (Na 2 B 4 O 7 .10H 2 O) are required to neutralize 25 mL each of 0.2 M H 2 SO 4 and H 3 PO 3 respectively are ?? que 2 An ore contains 2.5% of the mineral argentite, Ag 2 S, by mass. How many grams of this ore would have to be processed in order to obtain 0.675g of pure solid silver, Ag? Que 3 in a titration, 60 mL of Phosphoric acid (H 3 PO 4 ) completely neutralizes 750 mL of Sodium carbonate solution containing 3.18 gm of Na 2 CO 3 in one litre of solution. Thus the molarity of Phosphoric acid is??

 


Que. 1


The milli moles of Borax (Na2B4O7.10H2O) are required to neutralize 25 mL each of 0.2 M H2SO4 and H3PO3 respectively are ??


que 2



An ore contains 2.5% of the mineral argentite, Ag2S, by mass. How many grams of this ore would have to be processed in order to obtain 0.675g of pure solid silver, Ag?


 


Que 3



in a titration, 60 mL of Phosphoric acid (H3PO4) completely neutralizes 750 mL of Sodium carbonate solution containing 3.18 gm of Na2CO3


 


in one litre of solution. Thus the molarity of Phosphoric acid is??




 

Grade:11

1 Answers

Aman Bansal
12 years ago

Dear Aman,

1)calculate milliequivalents of h3po2 by multiplying volume and normality
No of milliequivalents = 100 * 0.2 = 20 milllieqs
millieqs of mg(oh)2 is 80 * 0.2 = 16millieqs
so 4 millieqs are left for nuetralization
thus we require 4 millieqs of koh
so 4 = vol * normality 
so 4 = volume * 0.25 
volume = 16ml
2)in ques 2 calculate the milli moles of h2so4 by multiplying molarity and volume ( in ml)
similarly for h3po3
then apply stoichiometry 
to the eqaution Na2B4O7 + 2HCl + 5H2O --> 2NaCl + 4H3BO3
millimoles of hcl = 25*0.2 = 5 millimoles
acc to equation 1mole of borax react with 2 moles of hcl
so 2.5 moles of borax react with 5 moles of hcl
similarly for h3po3
3)let the mass of ore be x
then 2.5 * x /100 = molecular wt of ag2s = 248
so x = 9920
so 9920 gram of ore yields 216 grams of silver ( ag2)
so 31 grams of ore with 0.675 gram of ag

Best Of luck

Plz Approve the answer...!!!!

Cracking IIT just got more exciting,It’s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and winexciting gifts by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple click here to download the toolbar….
So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert